OpenStudy (anonymous):

solution of the equation x^2-12x=8?

OpenStudy (anonymous):

2 root11+6 -2 root11+6

OpenStudy (bradely):

x^2-12x=8 Add 36 add both sides x^2-12x+36=8+36 (x-6)^2=44 x-6=+or- sqrt(44) x=6+or- sqrt(44) =6+or- 2sqrt(11) Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

\[ x^2-12x=8\\ (x-6)^2=8+36\\ (x-6)^2=44\\ x-6=\pm\sqrt{44}\\ x=6\pm2\sqrt{11}\]

OpenStudy (anonymous):

The equation for quadratic formulas to be solved is -b plus or minus root b2-4ac over 2a